OpenStudy (anonymous):

CALCULUS II - Average Values f(x) = 6x/((1+x^2)^2), [0,2] a) Find the average value f ave of f on the given interval. b) Find c such that f ave = f(c). There should be two answers. Round your answers to three decimal places.

OpenStudy (anonymous):

I finished A. A) 6/5 B) 6x/ ((1+x^2)^2) = 6/5

OpenStudy (rational):

A looks good

OpenStudy (rational):

you may use wolfram for part B http://www.wolframalpha.com/input/?i=solve+6%2F5+%3D++6x%2F%28%281%2Bx^2%29^2%29%2C+0%3Cx%3C2

OpenStudy (anonymous):

f(0)=6(0)/(1+(0)(0))^5=0 f(2)=6(2)/{1+2^2}^2=12/25 Av={0+(12/25)}/2=?

OpenStudy (anonymous):

Correction f(0)=6*0/{ 1+0^2}^2=0